Two cars leave City A at the same time to travel to city B. One car travels at 72 km/h and other at 78 km/h. If the slower car arrives 20 min. after the faster car, how far apart are City A and City B?
Answer by stanbon(48558)

Two cars leave City A at the same time to travel to city B. One car travels at 72 km/h and other at 78 km/h. If the slower car arrives 20 min. after the faster car, how far apart are City A and City B?
---------------
1st Car DATA:
Rate = 72 km/h ; Time = x hrs ; distance = 72x km
---------
2nd Car DATA:
Rate = 78 km/h ; Time = (x-(1/3)) hrs ; distance = 78(x-(1/3)) km
-----------
EQUATION:
distance = distance
72x = 78(x - (1/3))
72x = 78x - 26
6x= 26
x = 13/3 hrs
---------
So, how far is it from A to B?
72 km/h *(x hr) = 72(13/3) = 24*13 = 312 km
================
Cheers,
Stan H.